Diffusion Morning Report 8/21/03

ted berg tberg at snf.stanford.edu
Thu Aug 21 07:47:34 PDT 2003


Good Morning All,
    Everything is looking pretty good this morning.  LTO is being pulled 
and cleaned as of now.  Nitride is running  and testing.
TEOS had Temp issues yesterday,  we could find no smoking gun.

Diffusion 	Date: 	8/20/2003 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-Dirty Ox 	Up 	  	 
Tylan 8-SiGe 	Up 	  	9.9/30
Tylan 9-Poly 	Up 	  	19.22/30
Tylan 10-Nitride 	yellow 	Process testing 	11./30
Tylan 11-BPSG-LTO 	Down 	Pull and clean 	0/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	9.3/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	Temperature issues 	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	Currently 660- B(c7) 
on line


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20030821/324c0fd6/attachment.html>


More information about the diffusion mailing list